# Sparse-PGD: A Unified Framework for Sparse Adversarial Perturbations Generation

### v3

@@ 5 Adversarial Training
Random Backward Function: Since the sparse gradient and the unprojected gradient as described in Section 3.1 induce different exploration-exploitation trade-offs, we randomly select one of them to generate adversarial perturbations for each mini-batch when using sPGD to generate adversarial perturbations. Compared with mixing these two backward functions together, as in sAA, random backward function does not introduce computational overhead.
@@
6 Experiments
In this section, we conduct extensive experiments to compare our attack methods with baselines in evaluating the robustness of various models against $l_{0}$ bounded and structured sparse perturbations. Besides the effectiveness with an abundant query budget, we also study the efficiency of our methods. Our results demonstrate that our sPGD performs best among white-box attacks. With limited iterations, sPGD achieves significantly better performance than existing methods. Therefore, sAA, consisting of the best white-box and black-box attacks, has the best attack success rate. sPGD, due to its efficiency, is utilized for adversarial training to obtain the best robust models against $l_{0}$ bounded and structured sparse adversarial perturbations. To further demonstrate the efficacy of our method, we evaluate the transferability of sPGD. The results show that sPGD has a high transfer success rate, making it applicable in practical scenarios. The adversarial examples generated by our methods are presented in Sec. 6.7. In addition, we conduct ablation studies for analysis in C. Implementation details are deferred to Appendix A.
6.1 Robustness against Unstructured Sparse Perturbations
Table 1: Robust accuracy of various models on different sparse attacks, where the sparsity level $\epsilon=20$. The models are trained on CIFAR-10. Note that we report results of Sparse-RS (RS) with fine-tuned hyperparameters, which outperforms its original version in Croce et al. [17]. CornerSearch (CS) is evaluated on $1000$ samples due to its high computational complexity.
@@ 6.1 Robustness against Unstructured Sparse Perturbations
First, we compare our proposed sPGD, including sPGDproj (sPGDp) and sPGDunproj (sPGDu) as defined in Section 3.2, and sAA with existing white-box and black-box attacks that generate $l_{0}$ bounded sparse perturbations. We evaluate different attack methods based on the models trained on CIFAR-10 [64] and report the robust accuracy with $\epsilon=20$ on the whole test set in Table 1. Additionally, the results on ImageNet100 [65] and a real-world traffic sign dataset GTSRB [66] are reported in Table 2. Note that the image sizes in GTSRB vary from $15\times 15$ to $250\times 250$. For convenience, we resize them to $224\times 224$ and use the same model architecture as in ImageNet-100. Furthermore, only the training set of GTSRB has annotations, we manually split the original training set into a test set containing $1000$ instances and a new training set containing the rest data. In Appendix B, we report more results on CIFAR-10 with $\epsilon=10$, $\epsilon=15$, and the results of models trained on CIFAR-100 [64] in Table 7, 8, 9, respectively, to demonstrate the efficacy of our methods.
@@
Based on the results in Table 1 and Table 2, we can find that SF attack, PGD0 attack and SAIF attack perform significantly worse than our methods for all the models studied. That is, our proposed sPGD always performs the best among white-box attacks. Among black-box attacks, CS attack can achieve competitive performance, but it runs dozens of times longer than our method does. Therefore, we focus on comparing our method with RS attack. For $l_{1}$ and $l_{2}$ models, our proposed sPGD significantly outperforms RS attack. By contrast, RS attack outperforms sPGD for $l_{\infty}$ and $l_{0}$ models. This gradient masking phenomenon is, in fact, prevalent across sparse attacks. Given sufficient iterations, RS outperforms all other existing white-box attacks for $l_{\infty}$ and $l_{0}$ models. Nevertheless, among white-box attacks, sPGD exhibits the least susceptibility to gradient masking and has the best performance. The occurrence of gradient masking in the context of $l_{0}$ bounded perturbations can be attributed to the non-convex nature of adversarial budgets. In practice, the perturbation updates often significantly deviate from the direction of the gradients because of the projection to the non-convex set. Similar to AA in the $l_{1}$, $l_{2}$ and $l_{\infty}$ cases, sAA consists of both white-box and black-box attacks for comprehensive robustness evaluation. It achieves the best performance in all cases in Table 1 and Table 2 by a considerable margin.
Table 2: Robust accuracy of various models on different sparse attacks. Our sAT model is trained with $\epsilon=1200$. (a) Results on ImageNet-100 [65]. (b) Results on GTSRB [66]. Note that the results of Sparse-RS (RS) with tuned hyperparameters are reported. All models are RN-34, and are evaluated on $500$ samples. The results of SparseFool (SF) and PGD0 are included due to their poor performance, and CornerSearch (CS) is not evaluated here due to its high computational complexity, i.e. nearly 1 week on one GPU for each run.
@@
In the case of $l_{0}$ adversarial training, the models are adversarially trained against sparse attacks. However, Figure 3 illustrates that the performance of RS attack drastically deteriorates with limited iterations (e.g., smaller than $100$), so RS is not suitable for adversarial training where we need to generate strong adversarial perturbations in limited number iterations. Empirical evidence suggests that employing RS with $20$ iterations for adversarial training, i.e., the same number of iterations as in other methods, yields trivial performance, so it is not included in Table 1 or Table 2 for comparison. In addition, models trained by PGD0-A and PGD0-T, which generate $l_{0}$ bounded perturbations, exhibit poor robustness to various attack methods, especially sAA. By contrast, the models trained by sAT and sTRADES show the strongest robustness, indicated by the comprehensive sAA method and all other attack methods. Compared with sAT, sTRADES achieves better performance in both robustness and accuracy. Finally, models trained with $l_{1}$ bounded perturbations are the most robust ones among existing non-$l_{0}$ training methods. It could be attributed to the fact that $l_{1}$ norm is the tightest convex relaxation of $l_{0}$ norm [70]. From a qualitative perspective, $l_{1}$ attacks also generate relatively sparse perturbations [20], which makes the corresponding model robust to sparse perturbations to some degree.
@@
6.2 Robustness against Structured Sparse Perturbations
Table 3: Robust accuracy of various models on different structured sparse attacks. (a) Results on CIFAR-10, all models are PRN-18. (b) Results on ImageNet-100, the test set contains $500$ samples, all models are RN-34. Note that the evaluated group sparsity levels are approximately equivalent to the $l_{0}$ sparsity levels used in Table 1 and 2, and the Sparse-RS (RS) used here is the proposed extended version.
@@ 6.2 Robustness against Structured Sparse Perturbations
Apart from unstructured sparse perturbations, we evaluate the effectiveness of our method in Algorithm 2 in generating structured sparse perturbations in this subsection. Given the paucity of available methods for comparison, we focus on comparing our method with the extended Sparse-RS (RS) and LOAP [16], which is a white-box approach generating adversarial patches, on CIFAR-10 and ImageNet-100. For a comprehensive evaluation, we include the results of different types of structured sparse perturbations, e.g., row and patches with different group sparsity level and different sizes. It should be noted that the unstructured $l_{0}$ norms of the structured perturbations studied here are approximately the same as the perturbations in Table 1 and 2. In addition, we evaluate the attacks on vanilla models, models trained with unstructured $l_{0}$ bounded perturbations, and those trained with the specific structured sparse perturbations.
The results in Table 3 indicate that the proposed sAA also achieves the state-of-the-art performance in generating structured sparse perturbations in almost all cases. Furthermore, the models trained with our methods exhibit the strongest robustness against structured sparse perturbations. As anticipated, models trained on specific structured sparse adversarial examples significantly outperform those trained on unstructured sparse adversarial examples. This highlights the inherent limitation of adversarially trained models in maintaining robustness against unseen types of perturbations during training.
6.3 Adversarial Watermarks
Table 4: Robust accuracy of vanilla model (RN-34) against adversarial watermarks with different $l_{\infty}$ adversarial budgets. The experiment is conducted on ImageNet-100.
@@ 6.3 Adversarial Watermarks
(b) $60\times 60$ letter “A”
To evaluate the performance of our method in generating adversarial watermarks, we conduct experiments with different patterns and different $\l_{\infty}$ adversarial budgets. From Table 4, we can find that when the size of the pattern is large, like $60\times 60$ letter “A” with $1747$ non-zero elements, our attack can still achieve decent performance under different $l_{\infty}$ adversarial budgets. In contrast, when we adopt a $18\times 18$ circle pattern with $208$ non-zero elements, the attack success rate abruptly declines as the $l_{\infty}$ adversarial budget decreases. Nevertheless, from the aspect of attack success rate, our approaches still outperform Sparse-RS by a large margin.
6.4 Comparison under Different Iteration Numbers and Different Sparsity Levels
@@
Figure 3: Comparison between sPGD and RS attack under different iteration numbers and different sparsity levels. (a) Different iteration number comparison on CIFAR-10, $\epsilon=20$. ResNet18 (std), PORT ($l_{\infty}$ and $l_{2}$) [44], $l_{1}$-APGD ($l_{1}$) [19] and sTRADES ($l_{0}$) are evaluated. (b) Different iteration number comparison on ImageNet-100, $\epsilon=200$. ResNet34 (std), Fast-EG-$l_{1}$ ($l_{1}$) [20] and sAT ($l_{0}$) are evaluated. In (a) and (b), the total iteration number ranges from $20$ and $10000$. For better visualization, the x-axis is in the log scale. (c) Different sparsity comparison on CIFAR-10. The evaluated models are the same as those in (a). The $\epsilon$ ranges from 00 and $50$. The number of total iterations is set to $10000$. Note that the results of sPGD and RS attack are shown in solid lines and dotted lines, respectively.
@@
As illustrated in Figure 3 (a) and (b), sTRADES has better performance than other robust models by a large margin in all iterations of both sPGD and RS attacks, which is consistent with the results in Table 12 and 9. For vanilla and other robust models, although the performances of both sPGD and RS attack get improved with more iterations, sPGD outperforms RS attack by a large margin when the iteration number is small (e.g. $<1000$ iterations), which makes it feasible for adversarial training. Similar to other black-box attacks, the performance of RS attack drastically deteriorates when the query budget is limited. In addition, our proposed gradient-based sPGD significantly outperforms RS on ImageNet-100, where the search space is much larger than that on CIFAR-10, i.e., higher image resolution and higher sparsity level $\epsilon$. This suggests that our approach is scalable and shows higher efficiency on high-resolution images. Furthermore, although the performance of RS does not converge even when the iteration number reaches $10000$, a larger query budget will make it computationally impractical. Following the setting in Croce et al. [17], we do not consider a larger query budget in our experiments, either.
@@
6.5 Efficiency of sPGD
@@
Figure 4: Distribution of the iteration numbers needed by sPGD (blue) and RS (orange) to successfully generate adversarial samples. The results are obtained from different models: (a) $l_{\infty}$ PORT [44], (b) $l_{1}$-APGD [19] and (c) our sPGD. The average iteration numbers (Avg.) and attack success rate (ASR), i.e., $1-$Robust Acc., are reported in the legend. For better visualization, we clip the minimum iteration number to $10$ and show the x- and y-axis in log scale.
Table 5: Runtime of different attacks on 1000 test instances with batch size 500. The sparsity level $\epsilon=20$. The evaluated model is sTRADES. The model is trained on CIFAR-10. The experiments are implemented on NVIDIA Tesla V100.
@@
To further showcase the efficiency of our approach, we first compare the distributions of the number of iterations needed by sPGD and RS to successfully generate adversarial samples. Figure 4 illustrates the distribution of the iteration numbers needed by sPGD and RS to successfully generate adversarial samples. For $l_{\infty}$ and $l_{1}$ robust models, our proposed sPGD consumes distinctly fewer iteration numbers to successfully generate an adversarial sample than RS, the strongest black-box attack in Table 1, while maintaining a high attack success rate. Similar to the observations in Figure 3, the model trained by sTRADES suffers from gradient masking. However, RS still requires a large query budget to successfully generate an adversarial sample. This further demonstrates the efficiency of sPGD, which makes adversarial training feasible.
Additionally, we compare the runtime of sPGD with other attacks in the same configuration as in Table 1. As shown in Table 5, the proposed sPGD shows the highest efficiency among various attacks. Although sAA consumes more time (approximately 2×2$\times$2 × sPGD + RS), it can provide a reliable evaluation against $l_{0}$ bounded perturbation.
6.6 Transferability of Adversarial Perturbations
Table 6: Transferability of RS and sPGD between vanilla VGG11 (V) and RN-18 (R) on CIFAR-10 with $\epsilon=20$. Attack success rate (ASR) is reported. The perturbations are generated on the source model (left), and are evaluated on the target model (right). Note that $\alpha$ and $\beta$ of sPGD are set to $0.75$.
[…]

### v4

@@ V Adversarial Training
Random Backward Function: Since the sparse gradient and the unprojected gradient as described in Section III-A induce different exploration-exploitation trade-offs, we randomly select one of them to generate adversarial perturbations for each mini-batch when using sPGD to generate adversarial perturbations. Compared with mixing these two backward functions together, as in sAA, random backward function does not introduce computational overhead.
@@
VI Experiments
In this section, we conduct extensive experiments to compare our attack methods with baselines in evaluating the robustness of various models against $l_{0}$ bounded and structured sparse perturbations. Besides the effectiveness with an abundant query budget, we also study the efficiency of our methods. Our results demonstrate that our sPGD performs best among white-box attacks. With limited iterations, sPGD achieves significantly better performance than existing methods. Therefore, sAA, consisting of the best white-box and black-box attacks, has the best attack success rate. sPGD, due to its efficiency, is utilized for adversarial training to obtain the best robust models against $l_{0}$ bounded and structured sparse adversarial perturbations. To further demonstrate the efficacy of our method, we evaluate the transferability of sPGD. The results show that sPGD has a high transfer success rate, making it applicable in practical scenarios. Moreover, to exhibit the versatility of our method, we evaluate sPGD in objection detection and segmentation tasks. The adversarial examples generated by our methods are presented in Sec. VI-H. In addition, we conduct ablation studies for analysis in Appendix C. Implementation details are deferred to Appendix A.
TABLE I: Robust accuracy of various models on different sparse attacks. The models are trained on CIFAR-10. Note that we report results of Sparse-RS (RS) with fine-tuned hyperparameters, which outperforms its original version in Croce et al. [17]. CornerSearch (CS) is evaluated on 1000 samples due to its high computational complexity.
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
Vanilla RN-18 93.9 1.6 0.0 25.3 2.1 12.0 0.0 0.0 0.0
$l_{\infty}$-adv. trained, $\epsilon=8/255$
GD PRN-18 87.4 30.5 12.2 61.1 36.0 51.3 17.1 24.3 11.3
PORT RN-18 84.6 30.8 15.2 62.1 31.4 52.1 17.4 23.0 13.0
DKL WRN-28 92.2 35.3 13.2 62.5 41.2 52.3 18.4 24.9 12.1
DM WRN-28 92.4 34.8 12.6 57.9 38.5 49.4 17.9 24.0 11.6
$l_{2}$-adv. trained, $\epsilon=0.5$
HAT PRN-18 90.6 38.9 23.5 65.3 35.4 60.2 19.0 18.6 16.6
PORT RN-18 89.8 36.8 20.6 64.3 30.6 59.7 16.0 15.5 13.8
DM WRN-28 95.2 48.5 27.7 68.2 47.5 70.9 25.0 26.7 22.1
FDA WRN-28 91.8 47.8 31.1 71.8 40.1 68.2 28.0 31.4 25.5
$l_{1}$-adv. trained, $\epsilon=12$
$l_{1}$-APGD PRN-18 80.7 41.3 36.5 70.3 50.5 62.3 30.4 31.3 29.0
Fast-EG-$l_{1}$ PRN-18 76.2 40.7 34.8 64.9 46.7 56.9 29.6 30.1 28.0
$l_{0}$-adv. trained, $\epsilon=15$
PGD0-A PRN-18 83.7 17.5 6.1 73.7 62.9 60.5 19.4 27.5 5.6
PGD0-T PRN-18 90.5 19.5 7.2 85.5 63.6 69.8 31.4 41.2 7.1
sAT PRN-18 80.9 46.0 37.6 77.1 74.1 72.3 71.2 70.3 37.6
sTRADES PRN-18 90.3 71.7 63.7 89.5 88.1 86.5 85.9 83.8 63.7
(a) $\epsilon=$ 15
@@ VI Experiments
(b) $\epsilon=$ 20
VI-A Robustness against Unstructured Sparse Perturbations
First, we compare our proposed sPGD, including sPGDproj (sPGDp) and sPGDunproj (sPGDu) as defined in Section III-B, and sAA with existing white-box and black-box attacks that generate $l_{0}$ bounded sparse perturbations. We evaluate different attack methods based on the models trained on CIFAR-10 [64] and report the robust accuracy with $\epsilon=15,~20$ on the whole test set in Table II(b). Additionally, the results on ImageNet-100 [65] and a real-world traffic sign dataset GTSRB [66] are reported in Table II. Note that the image sizes in GTSRB vary from $15\times 15$ to $250\times 250$. For convenience, we resize them to $224\times 224$ and use the same model architecture as in ImageNet-100. Furthermore, only the training set of GTSRB has annotations, we manually split the original training set into a test set containing $1000$ instances and a new training set containing the rest data. In Appendix B, we report more results on CIFAR-10 and CIFAR-100 [64] in Table VIII, to demonstrate the efficacy of our methods.
@@ VI-A Robustness against Unstructured Sparse Perturbations
Based on the results in Table II(b) and Table II, we can find that SF attack, PGD0 attack and SAIF attack perform significantly worse than our methods for all the models studied. That is, our proposed sPGD always performs the best among white-box attacks. Among black-box attacks, CS attack can achieve competitive performance, but it runs dozens of times longer than our method does. Therefore, we focus on comparing our method with RS attack. For $l_{1}$ and $l_{2}$ models, our proposed sPGD significantly outperforms RS attack. By contrast, RS attack outperforms sPGD for $l_{\infty}$ and $l_{0}$ models. This gradient masking phenomenon is, in fact, prevalent across sparse attacks. Given sufficient iterations, RS outperforms all other existing white-box attacks for $l_{\infty}$ and $l_{0}$ models. Nevertheless, among white-box attacks, sPGD exhibits the least susceptibility to gradient masking and has the best performance. The occurrence of gradient masking in the context of $l_{0}$ bounded perturbations can be attributed to the non-convex nature of adversarial budgets. In practice, the perturbation updates often significantly deviate from the direction of the gradients because of the projection to the non-convex set. Similar to AA in the $l_{1}$, $l_{2}$ and $l_{\infty}$ cases, sAA consists of both white-box and black-box attacks for comprehensive robustness evaluation. It achieves the best performance in all cases in Table II(b) and Table II by a considerable margin.
TABLE II: Robust accuracy of various models on different sparse attacks. Our sAT model is trained with $\epsilon=1200$. (a) Results on ImageNet-100 [65]. (b) Results on GTSRB [66]. Note that the results of Sparse-RS (RS) with tuned hyperparameters are reported. All models are RN-34, and are evaluated on $500$ samples. The results of SparseFool (SF) and PGD0 are included due to their poor performance, and CornerSearch (CS) is not evaluated here due to its high computational complexity, i.e. nearly 1 week on one GPU for each run.
@@
In the case of $l_{0}$ adversarial training, the models are adversarially trained against sparse attacks. However, Figure 3 illustrates that the performance of RS attack on standard, $l_{1}$-, $l_{2}$- and $l_{\infty}$-trained models drastically deteriorates with limited iterations (e.g., smaller than $100$), so RS is not suitable for adversarial training where we need to generate strong adversarial perturbations in limited number iterations. Empirical evidence suggests that employing RS with $20$ iterations for adversarial training, i.e., the same number of iterations as in other methods, yields trivial performance, so it is not included in Table II(b) or Table II for comparison. In addition, models trained by PGD0-A and PGD0-T, which generate $l_{0}$ bounded perturbations, exhibit poor robustness to various attack methods, especially sAA. By contrast, the models trained by sAT and sTRADES show the strongest robustness, indicated by the comprehensive sAA method and all other attack methods. Compared with sAT, sTRADES achieves better performance in both robustness and accuracy. Finally, models trained with $l_{1}$ bounded perturbations are the most robust ones among existing non-$l_{0}$ training methods. It could be attributed to the fact that $l_{1}$ norm is the tightest convex relaxation of $l_{0}$ norm [69]. From a qualitative perspective, $l_{1}$ attacks also generate relatively sparse perturbations [20], which makes the corresponding model robust to sparse perturbations to some degree.
@@
VI-B Robustness against Structured Sparse Perturbations
TABLE III: Robust accuracy of various models on different structured sparse attacks. (a) Results on CIFAR-10, all models are PRN-18. (b) Results on ImageNet-100, the test set contains $500$ samples, all models are RN-34. Note that the evaluated group sparsity levels are approximately equivalent to the $l_{0}$ sparsity levels used in Table II(b) and II, and the Sparse-RS (RS) used here is the proposed extended version.
@@ VI-B Robustness against Structured Sparse Perturbations
Apart from unstructured sparse perturbations, we evaluate the effectiveness of our method in Algorithm 2 in generating structured sparse perturbations in this subsection. Given the paucity of available methods for comparison, we focus on comparing our method with the extended Sparse-RS (RS) and LOAP [16], which is a white-box approach generating adversarial patches, on CIFAR-10 and ImageNet-100. For a comprehensive evaluation, we include the results of different types of structured sparse perturbations, e.g., row and patches with different group sparsity level and different sizes. It should be noted that the unstructured $l_{0}$ norms of the structured perturbations studied here are approximately the same as the perturbations in Table II(b) and II. In addition, we evaluate the attacks on vanilla models, models trained with unstructured $l_{0}$ bounded perturbations, and those trained with the specific structured sparse perturbations.
The results in Table III indicate that the proposed sAA also achieves the state-of-the-art performance in generating structured sparse perturbations in almost all cases. Furthermore, the models trained with our methods exhibit the strongest robustness against structured sparse perturbations. As anticipated, models trained on specific structured sparse adversarial examples significantly outperform those trained on unstructured sparse adversarial examples. This highlights the inherent limitation of adversarially trained models in maintaining robustness against unseen types of perturbations during training.
VI-C Adversarial Watermarks
TABLE IV: Robust accuracy of vanilla model (RN-34) against adversarial watermarks with different $l_{\infty}$ adversarial budgets. The experiment is conducted on ImageNet-100.
@@ VI-C Adversarial Watermarks
(b) $\times60$ letter “A”
[…]
VI-D Comparison under Different Iteration Numbers and Different Sparsity Levels
@@
Figure 3: Comparison between sPGD and RS attack under different iteration numbers and different sparsity levels. (a) Different iteration number comparison on CIFAR-10, $\epsilon=20$. ResNet18 (std), PORT ($l_{\infty}$ and $l_{2}$) [44], $l_{1}$-APGD ($l_{1}$) [19] and sTRADES ($l_{0}$) are evaluated. (b) Different iteration number comparison on ImageNet-100, $\epsilon=200$. ResNet34 (std), Fast-EG-$l_{1}$ ($l_{1}$) [20] and sAT ($l_{0}$) are evaluated. In (a) and (b), the total iteration number ranges from 2020 and $10000$. For better visualization, the x-axis is in the log scale. (c) Different sparsity comparison on CIFAR-10. The evaluated models are the same as those in (a). The $\epsilon$ ranges from 0 and $50$. The number of total iterations is set to $10000$. Note that the results of sPGD and RS attack are shown in solid lines and dotted lines, respectively.
@@
As illustrated in Figure 3 (a) and (b), sTRADES has better performance than other robust models by a large margin in all iterations of both sPGD and RS attacks, which is consistent with the results in Table II(b)$I$ and VIII. For vanilla and other robust models, although the performances of both sPGD and RS attack get improved with more iterations, sPGD significantly outperforms RS attack when the iteration number is small (e.g. <1000<1000 iterations), which makes it feasible for adversarial training. Similar to other black-box attacks, the performance of RS attack drastically deteriorates when the query budget is limited. In addition, our proposed gradient-based sPGD significantly outperforms RS on ImageNet-100, where the search space is much larger than that on CIFAR-10, i.e., higher image resolution and higher sparsity level $\epsilon$. This suggests that our approach is scalable and shows higher efficiency on high-resolution images. Furthermore, although the performance of RS does not converge even when the iteration number reaches $10000$, a larger query budget will make it computationally impractical. Following the setting in Croce et al. [17], we do not consider a larger query budget in our experiments, either.
@@
VI-E Efficiency of sPGD
To further showcase the efficiency of our approach, we first compare the distributions of the number of iterations needed by sPGD and RS to successfully generate adversarial samples. Figure 4 illustrates the distribution of the iteration numbers needed by sPGD and RS to successfully generate adversarial samples. For $l_{\infty}$ and $l_{1}$ robust models, our proposed sPGD consumes distinctly fewer iteration numbers to successfully generate an adversarial sample than RS, the strongest black-box attack in Table II(b), while maintaining a high attack success rate. Similar to the observations in Figure 3, the model trained by sTRADES suffers from gradient masking. However, RS still requires a large query budget to successfully generate an adversarial sample. This further demonstrates the efficiency of sPGD, which makes adversarial training feasible.
@@ VI-E Efficiency of sPGD
Figure 4: Distribution of the iteration numbers needed by sPGD (blue) and RS (orange) to successfully generate adversarial samples. The results are obtained from different models: (a) $l_{\infty}$ PORT [44], (b) $l_{1}$-APGD [19] and (c) our sPGD. The average iteration numbers (Avg.) and attack success rate (ASR), i.e., 11-Robust Acc., are reported in the legend. For better visualization, we clip the minimum iteration number to $10$ and show the x- and y-axis in log scale.
Additionally, we compare the runtime of sPGD with other attacks in the same configuration as in Table II(b). As shown in Table V, the proposed sPGD shows the highest efficiency among various attacks. Although sAA consumes more time (approximately $2\times$ sPGD + RS), it can provide a reliable evaluation against $l_{0}$ bounded perturbation.
TABLE V: Runtime of different attacks on 1000 test instances with batch size 500. The sparsity level $\epsilon=20$. The evaluated model is sTRADES. The model is trained on CIFAR-10. The experiments are implemented on NVIDIA Tesla V100.
@@
VI-F Transferability of Adversarial Perturbations
TABLE VI: Transferability of RS and sPGD between VGG11 (V) and ResNet18 (R) on CIFAR-10 with $\epsilon=$ 20. Attack success rate (ASR) is reported. The perturbations are generated on the source model (left), and are evaluated on the target model (right). Note that vanilla models are evaluated, and $\alpha$ and $\beta$ of sPGD are set to 0.75.
$\rightarrowV V\rightarrowR$ $\rightarrowR R\rightarrowV$